# Two transitions in complex eigenvalue statistics: Hermiticity and integrability breaking

### v1

@@ Abstract
Open quantum systems have complex energy eigenvalues which are expected to follow non-Hermitian random matrix statistics when chaotic, or 2-dimensional (2d) Poisson statistics when integrable. We investigate the spectral properties of a many-body quantum spin chain, the Hermitian XXZ Heisenberg model with imaginary disorder. Its rich complex eigenvalue statistics is found to separately break both Hermiticity and integrability at different scales of the disorder strength. With no disorder, the system is integrable and Hermitian, with spectral statistics corresponding to 1d Poisson. At very small disorder, we find a transition from 1d Poisson statistics to an effective $D$-dimensional Poisson point process, showing Hermiticity breaking. At intermediate disorder we find integrability breaking, and the statistics agrees with that of non-Hermitian complex symmetric random matrices in class AI. For large disorder, we recover the expected 2d Poisson statistics. Our analysis uses numerically generated nearest and next-to-nearest neighbour spacing distributions of an effective 2d Coulomb gas description at inverse temperature $\beta$, fitting them to the spin chain data. We confirm such an effective description of random matrices in class AI and AII up to next-to-nearest neighbour spacings.
@@ I Introduction
Quantum systems are in general not perfectly isolated, but rather interact with their surrounding environment. When the timescales of the environment are much shorter than the system’s ones and under the Born approximation, the reduced system’s dynamics can be described by the Lindblad master equation [Breuer2007Theory], or, effectively, by a Schrödinger equation with a non-Hermitian Hamiltonian [MingantiPRA2019, NaghilooNatPhys2019, RoccatiOSID2022]. In both cases, the operator governing the dynamics is non-Hermitian, and thus generically has a complex spectrum. Grobe, Haake and Sommers (GHS) [grobe1988quantum] conjectured that, for Markovian, open dissipative quantum systems, spectral properties can be used to distinguish between chaos and integrability, in a way similar to the Hermitian case. In particular, open quantum systems that are chaotic should display non-Hermitian random matrix statistics, while integrable systems with complex spectra should follow the statistics of a Poisson point process in two dimensions (2d Poisson). This conjecture was tested early on for the discrete quantum maps of periodically kicked tops with damping [grobe1988quantum] and in random neural networks [sompolinsky1988chaos], see below for further recent examples. In both cases, the nearest-neighbour spacing distribution was found to agree with 2d Poisson in the integrable case, and with that of the complex Ginibre ensemble [ginibre1965] in the fully chaotic regime.
@@
The model considered in this work, introduced in Ref. [roccati2024diagnosing], is peculiar as the disorder is purely imaginary, breaking at the same time both integrability and Hermiticity. This model is motivated by a recent experimental work [lapp2019engineering], showing that local tunable loss terms can be engineered in synthetic lattices. Moreover, the model allows us to investigate whether the breaking of Hermiticity happens on a different scale than the transition between chaotic and integrable behaviour. We quantitatively characterize the crossovers by comparing the nearest-neighbour (NN) level statistics of complex eigenvalues to that of a static Coulomb gas in 2d. In order to go beyond the smallest local scale, we will also use next-to-nearest neighbour (NNN) spacing distributions. The 2d Coulomb gas (2dCG) allows us to reproduce the 2d Poisson, and approximate the random matrix class AI as a function of the Coulomb gas inverse temperature $\beta$ [akemann2022spacing], thus enabling to capture a transition from integrable ($\beta=0)$ to chaotic ($\beta=1.4$) behavior, respectively. The breaking of Hermiticity is captured through a $D$-dimensional Poisson point process, with $D$ transitioning from a value of 1 for a purely Hermitian spectra to 2 for purely non-Hermitian systems.
The article is organised as follows. The next Section II describes our model, the XXZ spin chain with purely imaginary disorder, which is Hermitian and integrable at vanishing disorder, and its symmetries. In Section III, we introduce the complex eigenvalue statistics for NN and NNN spacing distributions we will use to characterize the model, namely an interpolating 2dCG ensemble for which an accurate surmise exists close to 2d Poisson, Poisson point process statistics (in varying dimension $D$), and the random matrix statistics of class AI, appropriate for the considered model. Section IV describes methods, in particular the unfolding procedure used in 2d and the tools to measure distances between two probability distributions. We briefly present complex spacing ratios introduced by [sa2020complex] as an alternative tool. The main Section V contains our results, the transition between random matrix and 2d Poisson statistics in $\beta$ of the 2dCG, and a second Hermiticity breaking crossover for very small disorder, between effectively 1d and 2d Poisson statistics. Our conclusions and open questions are presented in Section VI. Appendix A summarises the other two distinct random matrix symmetry, classes A and AII, including NN and NNN spacings.
@@ II.1 Definition of the model
describing site-dependent, local, dissipative disorder, because of the imaginary $i$ in (1). The site-dependent losses $\gamma_{a}$ can be experimentally implemented, as demonstrated in a synthetic lattice [lapp2019engineering]. Here, they are sampled from a uniform distribution $\mathcal{U}(0,\gamma)$, where $\gamma$ represents the ‘strength’ of the disorder. For each value of $\gamma$, we generate spectral data from a number of realizations of the model, each realization having a different randomly chosen set of local disorder terms, $\{\gamma_{a}\}_{a=1}^{L}$. We consider the case of periodic boundary conditions, as depicted in Figure 1.
Note that in this model, all the disorder and its implications to the spectral statistics come from the openness of the system—the spin chain Hamiltonian (2) is always fixed and is integrable according to the Bethe Ansatz [XXZ_bethe_ansatz]. For the study of spectral statistics of non-Hermitian models with real disorder see e.g. [hamazaki2019non, huang2020anderson, roccati2024diagnosing]. In [roccati2024diagnosing], it was shown that increasing the range $\gamma$, the statistics of the singular values exhibit a crossover from chaotic to integrable. The complex eigenvalue spacing statistics tell a similar story, as we show below, together with a richer behavior coming from Hermiticity breaking.
This model preserves the total spin in the $z$ direction, $\mathcal{S}_{z}=\sum_{a=1}^{L}S_{a}^{z}$. We thus work in a specific total $\mathcal{S}_{z}$ magnetization sector, usually the largest one with $\mathcal{S}_{z}=0$ which occurs for even $L$ and is of dimension $\binom{L}{L/2}$. At $\gamma=0$, additional symmetries need to be taken into account to avoid exact degeneracies in the spectrum (see e.g. [joel2013introduction]). Specifically, each $\mathcal{S}_{z}$ sector then splits into two: one sector with states of even Parity and another with states of odd Parity.
@@ II.2 Pseudo-Hermitian complex spectrum
which acts as $SS^{x}_{a}S^{-1}=+S^{x}_{a}$, $SS^{y}_{a}S^{-1}=-S^{y}_{a}$ and $SS^{z}_{a}S^{-1}=-S^{z}_{a}$ for any site $a$. In $H_{\textrm{XXZ}}$, the spin operators appear in pairs so the sign change is cancelled. In turn, the $S^{z}_{a}$ operators appear once in each term of $\Gamma$, leading to $S\Gamma S^{-1}=-\Gamma$, as needed to fulfill the symmetry (4).
III Spectral statistics of complex eigenvalues

### v2

@@ Abstract
Open quantum systems have complex energy eigenvalues which are expected to follow non-Hermitian random matrix statistics when chaotic, or 2-dimensional (2d) Poisson statistics when integrable. We investigate the spectral properties of a many-body quantum spin chain, the Hermitian XXZ Heisenberg model with imaginary disorder. Its rich complex eigenvalue statistics is found to separately break both Hermiticity and integrability at different scales of the disorder strength. With no disorder, the system is integrable and Hermitian, with spectral statistics corresponding to the 1d Poisson point process. At very small disorder, we find a transition from 1d Poisson statistics to an effective $D$-dimensional Poisson point process, showing Hermiticity breaking. At intermediate disorder we find integrability breaking, as infered from the statistics matching that of non-Hermitian complex symmetric random matrices in class AI. For large disorder, as the spins align, we recover the expected integrability (now in the non-Hermitian setup), indicated by 2d Poisson statistics. These conclusions are based on fitting the spin chain data of numerically generated nearest and next-to-nearest neighbour spacing distributions to an effective 2d Coulomb gas description at inverse temperature $\beta$. We confirm such an effective description of random matrices also applies in class AI and AII up to next-to-nearest neighbour spacings.
@@ I Introduction
Quantum systems are in general not perfectly isolated, but rather interact with their surrounding environment. When the timescales of the environment are much shorter than the system’s ones and under the Born approximation, the reduced system’s dynamics can be effectively described by a Schrödinger equation with a non-Hermitian Hamiltonian [MingantiPRA2019, NaghilooNatPhys2019, RoccatiOSID2022] and thus generically has a complex spectrum. Grobe, Haake and Sommers (GHS) [grobe1988quantum] conjectured that, for Markovian, open quantum systems, spectral properties can be used to distinguish between chaos and integrability, in a way similar to the Hermitian case. In particular, open quantum systems that are chaotic should display non-Hermitian random matrix statistics, while integrable systems with complex spectra should follow the statistics of a Poisson point process in two dimensions (2d Poisson). This conjecture was tested early on for the discrete quantum map of periodically kicked tops with damping [grobe1988quantum], and in random neural networks [sompolinsky1988chaos], see below for further recent examples. In both cases, the nearest-neighbour spacing distribution was found to agree with 2d Poisson in the integrable case, and with that of the complex Ginibre ensemble [ginibre1965] in the fully chaotic regime.
@@
The model considered in this work, introduced in Ref. [roccati2024diagnosing], has the peculiarity to have a purely imaginary disorder, breaking at the same time both integrability and Hermiticity. This model is motivated by a recent experimental work [lapp2019engineering], showing that local tunable loss terms can be engineered in synthetic lattices. Moreover, the model allows us to investigate whether the breaking of Hermiticity happens on a different scale than the transition between chaotic and integrable behaviour. We quantitatively characterize the crossovers by comparing the nearest-neighbour (NN) level statistics of complex eigenvalues to that of a static Coulomb gas in 2d. In order to go beyond the smallest local scale, we will also use next-to-nearest neighbour (NNN) spacing distributions. The 2d Coulomb gas (2dCG) allows us to reproduce the 2d Poisson, and approximate the random matrix class AI as a function of the Coulomb gas inverse temperature $\beta$ [akemann2022spacing], thus enabling to capture a transition from integrable ($\beta=0)$ to chaotic ($\beta=1.4$) behavior, respectively. The breaking of Hermiticity is captured through a $D$-dimensional Poisson point process, with $D$ transitioning from a value of 1 for a purely Hermitian spectrum to 2 for strongly non-Hermitian systems.
The article is organised as follows. The next Section II describes our model, the XXZ spin chain with purely imaginary disorder, which is Hermitian and integrable at vanishing disorder, and its symmetries. The main Section III contains our results, namely the transition between random matrix and 2d Poisson statistics in $\beta$ of the 2dCG, and, for very small disorder, the transition from 1d to 2d Poisson statistics, indicating a Hermiticity breaking crossover. In Section IV, we introduce the corresponding complex eigenvalue statistics for NN and NNN spacing distributions used to characterize the model, namely first, an interpolating 2dCG ensemble for which an accurate surmise exists close to 2d Poisson, second, Poisson point process statistics (in varying dimension $D$), and third, the random matrix statistics of class AI, appropriate for the considered model. Section V describes our methods, in particular the unfolding procedure used in 2d and the tools to measure distances between two probability distributions. For completeness, we briefly present complex spacing ratios introduced by [sa2020complex] as an alternative tool. Our conclusions and open questions are presented in Section VI. Appendix A summarises the other two distinct random matrix symmetry classes A and AII, including NN and NNN spacings.
@@ II.1 Definition of the model
representing site-dependent, local, dissipative disorder, because of the imaginary $i$ in (1). The site-dependent losses $\gamma_{a}$ can be experimentally implemented, as demonstrated in a synthetic lattice [lapp2019engineering]. Here, they are sampled from a uniform distribution $\mathcal{U}(0,\gamma)$, where $\gamma$ represents the ‘strength’ of the disorder. For each value of $\gamma$, we generate spectral data from a number of realizations of the model, each realization having a different randomly chosen set of local disorder terms, $\{\gamma_{a}\}_{a=1}^{L}$. We consider the case of periodic boundary conditions, as depicted in Figure 1.
Note that in this model, all the disorder and its implications to the spectral statistics come from the openness of the system—the spin chain Hamiltonian (2) is always fixed and is integrable according to the Bethe Ansatz [XXZ_bethe_ansatz]. For the study of spectral statistics of non-Hermitian models with real disorder see e.g. [hamazaki2019non, huang2020anderson, roccati2024diagnosing]. In [roccati2024diagnosing], it was shown that increasing the range $\gamma$, the statistics of the singular values exhibit a crossover from chaotic to integrable. A similar observation is made here from the eigenvalue spacing statistics. We complete the story by extending the analysis to very small disorder, which reveals a richer behavior coming from Hermiticity breaking. For some aspects of thermalization and transport in this model, see [Mahoney:2024sgu]. This model preserves the total spin in the $z$ direction, $\mathcal{S}_{z}=\sum_{a=1}^{L}S_{a}^{z}$. We thus work in a specific total $\mathcal{S}_{z}$ magnetization sector, usually the largest one with $\mathcal{S}_{z}=0$ which occurs for even $L$ and is of dimension $\binom{L}{L/2}$. Without disorder ($\gamma=0$), additional symmetries need to be taken into account to avoid exact degeneracies in the spectrum (see e.g. [joel2013introduction]). Specifically, each $\mathcal{S}_{z}$ sector then splits into two: one sector with states of even Parity and another with states of odd Parity.
@@ II.2 Pseudo-Hermitian complex spectrum
which acts as $SS^{x}_{a}S^{-1}=+S^{x}_{a}$, $SS^{y}_{a}S^{-1}=-S^{y}_{a}$ and $SS^{z}_{a}S^{-1}=-S^{z}_{a}$ for any site $a$. In $H_{\textrm{XXZ}}$, the spin operators appear in pairs and so the sign change is cancelled. In turn, the $S^{z}_{a}$ operators appear once in each term of $\Gamma$, leading to $S\Gamma S^{-1}=-\Gamma$, as needed to fulfill the symmetry (4).
III Results for the spin chain
We now characterize the spectral statistics of the spin chain with dissipative disorder described by the non-Hermitian Hamiltonian (1). We do so by comparing the distributions of its eigenvalue spacings with that of the 2d spectral statistics of the reference models, detailed in Sec. IV, using the methods presented in Sec. V. We focus on the NN and NNN spacing distributions for values of the disorder strength from $\gamma=0-20$, fitting the numerical distribution to either the 2dCG distributions and finding the corresponding $\beta$, or to a $D$-dimensional Poisson point process with $1\leq D\leq 2$. For $L=16$, we present NN, NNN, and complex spacing ratio data for over 450 disorder realizations, while for $L=12,14$, we show this data for 1000 disorder realizations. To complete the picture, a qualitative discussion of the complex spacing ratios [sa2020complex] is included at the end of the section.
We expect the spectral statistics of our model to behave similarly to those of a XXZ spin-chain Hamiltonian with real disorder, which exhibits a crossover of the spectral statistics from chaos to integrability (for non-zero disorder strength) as the disorder strength is increased. In such a case, the NN spectral statistics changes from random matrix statistics approximated by the Wigner surmise $\propto se^{-as^{2}}$ for the GOE in this model to 1d Poisson, $\propto e^{-s}$, $s$ being the spacing between NN eigenergies. In 1d, there is universal level repulsion between the eigenvalues only in the first, chaotic regime, while in the second, 1d Poisson displays uncorrelated eigenvalues—a universal signature of generic integrable (deterministic) systems.
Figure 2: (a) Spacing distributions of the Poisson process, with dimension $1\leq D\leq 2$ varying in steps $0.1$ (red–grey). Notice that in both cases the maximum moves from right to left when decreasing dimension $D$. (b) The numerically generated NN and NNN spacing distributions of the 2dCG for $\beta=0-2$, increasing in steps $0.2$ (grey–blue). This also includes analytic formulas for 2d Poisson with $\beta=0$, (9) and (10) with the lowest maxima, and random matrix class A with $\beta=2$, (20) and (LABEL:ANNN), with highest maxima. Note that the grey curves for $D=2$ in (a) are the same as the grey curves for $\beta=0$ in (b) since the two processes then coincide.
In the non-Hermitian (2d) case with imaginary disorder, where the eigenvalues lie in the complex plane, we expect the spectral statistics to exhibit a crossover between random matrix statistics of class AI in our case to that of the 2d Poisson point process—which has a linear repulsion from the area measure, see (9) at $D=2$. We follow the proposal [akemann2019universal] and use the spectral statistics of a 2d Coulomb gas with some value for $\beta\geq 0$, which includes 2d Poisson at $\beta=0$ as a special case and approximates the AI class for $\beta\approx 1.3-1.4$ [akemann2022spacing]. Fitting $\beta$ thus allows us to capture this crossover. However, the transition has more structure than initially expected, showing two different crossovers, as we summarise below and detail in subsequent subsections.
On a technical level, the spin chain data need to be drawn away from the edges of the eigenvalue distributions, where the statistics may be different or not even universal, as discussed at the beginning of Section IV. To do so, we focus on the bulk given by the denser inner part of the spectrum. We systematically select only that part of the spectrum for which the eigenvalue density is not less than some fraction of the maximal density. For $L=16$ spins with $\gamma=0.75-20$, we use $1/2$ of the maximal density; for $L=14$ and $L=12$, we use a $1/3$ of the maximal density as a cutoff for the eigenvalues to be included in the analysis. For $L=16$ spins with $\gamma<0.75$ we choose a cutoff equal or smaller to $1/4$ the maximal density because, as we discuss below, the spectrum is very peaked around a line parallel to the real line.
The 2d spectral statistics of the $D$-dimensional process and 2dCG, detailed in Section IV, are presented in Fig. 2. For our analysis, the relevant features are as follows. The NN distribution of the 2dCG has its leftmost and lowest maximum at $\beta=0$, see Fig. 2(b) left (grey curve). Increasing $\beta$ leads to a monotonous increase of the maximum in height and in position to the right (grey to blue). The random matrix class AI expected at maximal chaos is the one with the lowest maximum of the three distinct random matrix classes (see Fig. A1), reached at approximately $\beta\approx 1.3-1.4$. As such, it is approached by the 2dCG monotonously from below, starting at $\beta=0$. What happens if the maximum of the NN spacing distribution of our spin chain data—which is normalised and has first moment of unity—is to the left of the 2d Poisson distribution (at $\beta=0$)? Obviously the 2dCG will no longer provide a good fit then. Only the $D$-dimensional Poisson distribution might provide a good fit then for $D<2$, see Fig. 2(a) left. There, it is shown that the maximum of the distribution goes from left to right as the effective dimension is increased.
Our findings can be summarised by Fig. 3 and described as follows. In the small disorder regime, there is a crossover between 1d Poisson at $\gamma=0$—where our XXZ spin chain is integrable— towards 2d Poisson, that is well described by an increase in effective dimension $21\leq D<$ 2. We reach up to $D\approx 1.7$ at $\gamma=0.25$ for both NN and NNN spacing distributions. Therefore, Hermiticity is broken first on a smaller scale. This is summarised in Fig. 3 left and described in detail in Subsection III.1.
For intermediate to large values of $\gamma=0.5-20$, we see a rise from $\beta\approx 0.6$ at $\gamma=0.5$ to $\beta\approx 1.3-1.4$ at $\gamma=2$ where class AI is located, to which we also compare. We see that integrability is broken as we approach the fully chaotic behaviour characterised by random matrix class AI. For larger $\gamma$, the disorder term (3) leads to an alignment of spins and thus again to a deterministic behaviour, approaching 2d Poisson with $\beta=0$ around $\gamma\approx 8$, see Fig. 3 right. This is discussed in Subsection III.2, including the approach to class AI with system size $L$ in Subsection III.2.1. The asymptotic approach to $\beta=0$ between $\gamma=8-20$ is discussed in Subsection III.3, where we also use the surmise (7) for the finer scaling in the NN distribution. To complete the picture, we show and discuss results for the complex spacing ratios at various values of $\gamma$ in Subsection III.4.
III.1 Hermiticity breaking at small disorder $\gamma$:
from 1d to 2d Poisson
Starting with $\gamma=0$ (left plots), the spectrum is real, the XXZ model integrable, and we obtain a very good agreement with 1d Poisson statistics for both NN and NNN as expected. Notice that at $\gamma=0$ the energies are doubly degenerate, which is why we only show spacings with $\mathcal{S}_{z}=0$ and parity $P=+1$ (for $L=16$, this sector is of dimension 6470). The plots for $\mathcal{S}_{z}=0$ with $P=-1$ (for $L=16$, this sector is of dimension 6400) look similar. For our smallest value of $\gamma$, we obtain a somewhat reasonable fit for the NN, with $D\approx 0.88$ consistent with $D=1$. In contrast, the distribution for the NNN does not yield a good fit at all, a feature that persists for $\gamma=0.1$. A possible explanation is that for these small values of $\gamma$ the spectrum is almost singular, being concentrated on a line parallel to the real line, with very few eigenvalues in the complex plane, see Fig. 5, and the unfolding procedure described in Section V.1 for a 2d complex spectrum may then not be completely suitable. For $\gamma=0$, we used the 1d version of the unfolding method described in V.1  [^2] We found that for 1d, $\Sigma=1.8\bar{s}$ gives good results..
$\gamma=0.5$               $\gamma=1$               $\gamma=2$                $\gamma=3$               $\gamma=4$               $\gamma=5$                $\gamma=10$
Eigenvalues
From $\gamma=0.1$ both NN and NNN start to deviate from $D=1$, and the best fits in $D$ reach values up to $D\approx 1.7$ at $\gamma=0.25$. Comparing NN and NNN leads to consistent values within a small margin. We thus believe that the fits to a one-parameter family of distributions are not a coincidence, but rather capture properties of an underlying effective Poisson interaction.
This leads to the interpretation that Hermiticity is broken first, while maintaining an (almost) integrable system, the characteristic of which is given by a Poisson distribution at intermediate dimension $21<D<$ 2. Increasing to the next higher value of $\gamma=0.5$, detailed in the next subsection III.2, the maximum of the NN crosses to the right of the 2d Poisson distribution (at $\beta=0$), which then fits the 2dCG. The fact that $D=2$ is not fully reached by increasing $\gamma$ here can be interpreted as a competition between Hermiticity and integrability breakings around the scale $\gamma=0.25-0.5$.
As an example for the crossover between one- and two-dimensional spectra, we show in Fig. 5 cuts through the spectral density perpendicular to the real axis at $x_{0}\approx 0$ as a function of imaginary part $y$. Noting the difference in scale, we see that between $\gamma=0.1$ and $0.2$ shown here, the spectrum slowly softens and starts spreading into the complex plane. The difference in abundance of eigenvalues with non-vanishing imaginary part may explain why the NNN distribution in Fig. 4 fits well only from a certain value of $\gamma=0.15$ onwards.
Table 1: List of best fit $\beta$ for the comparison of the NN spacing and NNN spacing of the spin chain with $L=16$ and $\gamma=0.5,0.75,1,2,\ldots 20$ to the 2dCG and the corresponding standard deviation $\sigma$ in units of $10^{-2}$. Notice that apart from $\gamma=0.5$ the values obtained for $\beta$ from NN and NNN agree or differ by $0.1$ only.
III.2 Integrability breaking at intermediate to large disorder $\gamma$: from 2d Poisson to class AI and back
In this section, we discuss the transition from random matrix statistics class AI to 2d Poisson statistics, witnessed for intermediate ($\gamma=0.5$) to large ($\gamma=20$) disorder dissipation strength. The approach towards fully chaotic statistics represented by random matrix class AI when increasing system size $L=12,14$ and 16 is included in subsection III.2.1 below.
The smallest value of disorder for which a meaningful fit with a 2dCG at $\beta\geq 0$ can be made is $\gamma=0.5$, when, $\beta=0.6\,(0.4)$ for NN (NNN), see Fig. 6 left. At this disorder strength, both Hermiticity and integrability appear to be broken. The system is in between the integrable case ($\beta=0$) and the fully chaotic case ($\beta\approx 1.3-1.4$, equivalent to random matrix symmetry class AI). The latter is then attained at $\gamma=2$, where the maximal value for $\beta$ (and height of the maximum) is reached, see Figure 3 right. We compare here the NN and NNN spacings with the 2dCG rather than AI. This is because the AI class is very well approximated by the 2dCG at these values of $\beta$, as shown in Fig. 11 and Table 4. Furthermore, random matrix statistics is approached in $\beta$ from below, see also the next subsection III.2.1, and the $\sigma$-value gives a measure of how close it is.
Increasing $\gamma$ further leads to a rapid decrease in the best fitted value for $\beta$, see also Table 1 below. From $\gamma\simeq 8-10$ onwards we consistently obtain $\beta=0$. This is because the disorder is so strong that the spins align with the disorder term and the statistics becomes again deterministic. The best fits for $\beta$ for NN and NNN and corresponding values, including standard deviations, are collected in Table 1. In this section we only fit to $\beta$ in the 2dCG where the step width is 0.1, but both NN and NNN distribution are available. The next section III.3 presents a much finer fit in $\beta$ for $\gamma\geq 8$, possible for the NN only thanks to the surmise (7), which is a very good approximation so close to $\beta=0$. The approach to zero can also be seen well in the summary Fig. 3 right. Similar findings for Hermitian Hamiltonians comparing to 1d Poisson and corresponding classical random matrices, using both NN and NNN spacings, were made very recently [kundu2024beyond].
$L=12$                $L=14$               $L=16$
NN
NNN
Figure 7: Comparison of the XXZ with dissipative order $\gamma=2$ in the $\mathcal{S}_{z}=0$ magnetization sector with increasing number of spins $L=12,14,16$ (from left to right, histograms) and the symmetry class AI of complex symmetric matrices (blue curve) of size $N=5000$. The NN distribution is given in the first row, whereas the NNN spacing is shown in the second row. Additionally, we give the standard deviation $\sigma$ between the two plotted distributions below each plot. We observe a decrease of the standard deviation for growing $L$, as expected.
III.2.1 Dependence on system size $L$
The main data presented so far are always for the longest spin chain, with $L=16$. This corresponds to a Hilbert space dimension of $12870$, that allows reaching a fully chaotic behaviour. Here, we also present data for system sizes $L=14$ and $L=12$, with respective dimensions of Hilbert space $3432$ and $924$, at the values $\gamma=2$ and $\gamma=20$. While for all three system sizes the 2d Poisson limit is reached fully at large $\gamma=20$, see Fig. 8, this is not the case for the limit towards random matrix statistics of class AI, see Fig. 7.
size $L$ 12 14 16
NN $\beta$ 0.9 1.2 1.3
$\sigma$ 1.3 0.6 0.7
NNN $\beta$ 0.8 1.1 1.3
$\sigma$ 3.5 1.3 1.0
Table 3: List of best fit $\beta$ for the comparison of the nearest-neighbour (NN) spacing of the spin chain with $L=16$ and $\gamma=8,\ldots 20$ to the NN 2dCG surmise (7) and the corresponding standard deviation $\sigma$ in units of $10^{-2}$.
$\boldsymbol{\gamma}$ 8 9 10 11 12 13 14 15 16 17 18 19 20
$\sigma$ 0.36 0.35 0.33 0.35 0.33 0.33 0.31 0.3 0.27 0.24 0.24 0.21 0.2
III.3 Asymptotic approach to 2d Poisson for large disorder $\gamma$
III.4 Comparison to complex spacing ratio discussion
The probability density of the complex spacing ratios, as seen in Figure 10, contains interesting patterns especially in the chaotic regime: the lower density around the origin and around $r=1$ are due to level repulsion. It is not likely for an eigenvalue and its NN to be close together, thus the density is lower at $r=0$; and an eigenvalue’s NN and NNN would most likely lie on either side of it, making the density lower at $r=1$; see also Figure 12 which compares the complex spacing ratio distribution of 2d Poisson with that of AI. The ‘bow and arrow’ feature seen for small $\gamma$ is a manifestation of an underlying structure of the complex eigenvalues. As observed and explained in Appendix D of [Li:2024uzg], this feature is a result of an eigenvalue’s NN and NNN being on the real line and as complex conjugate to it, or vice versa. Section II.2 showed that the Hamiltonian is pseudo-Hermitian and thus its spectrum is symmetric to reflection across a line parallel to the real axis. In the complex spacing ratios, this symmetry manifests itself as reflection symmetry across the real axis. Additionally, the excess of ratios along the real line is a result of real eigenvalues having real NN and NNN.
IV Spectral statistics of complex eigenvalues